OpenStudy (raden):

Find all pear (x,y) where x,y integer number and x not equal y, such that : (x^2+2y^2)/(2x^2+y^2) is a perfect square number ?

OpenStudy (anonymous):

So my brain has stopped working or this is trivial !!! \[\frac{x^2+2y^2}{2x^2+y^2}=n^2\]\[x^2+2y^2=(2x^2+y^2)n^2=2x^2n^2+y^2n^2\]\[x^2-2x^2n^2=y^2n^2-2y^2\]\[x^2(1-2n^2)=y^2(n^2-2)\]for \(n>1\) LHS is negative and RHS is positive...so \(n=1\) and \(y=\pm x\)

OpenStudy (anonymous):

@mukushla from this how can we find all possible pairs??

OpenStudy (anonymous):

well i proved that \(x=\pm y\) so all pairs are in the form \((x,y)=(\pm a,\pm a)\)

OpenStudy (experimentx):

x not equal y ,,, so (+a, -a) ..lol

OpenStudy (anonymous):

lol

OpenStudy (experimentx):

when i comes to Diophantine eqn ... i haven't found anyone better than you man!!

OpenStudy (anonymous):

@RadEn what do u think?

OpenStudy (raden):

so, there are infinity pairs, isn't mukushlah?

OpenStudy (anonymous):

yeah solutions are infinite

OpenStudy (raden):

Ok, thank u... sorry, my conection very low... i come late see it's answer :(

OpenStudy (anonymous):

:)

OpenStudy (raden):

wait, dont run every where.... i have one more, iwll repost soon ... :D

OpenStudy (anonymous):

ok @experimentX Oh santosh i see ur last reply now...that means a lot from u man...thank u

OpenStudy (experimentx):

no probs man!!